OpenStudy (rayford):

What are the possible rational zeros of f(x) = 2x3 – 4x2 – 7x + 14 ?

OpenStudy (anonymous):

set the function equal to zero and solve for x

OpenStudy (rayford):

huh? may you walk me through it?

OpenStudy (anonymous):

2x^3 – 4x^2 – 7x + 14 = 0

OpenStudy (rayford):

that's not a walk through :(...

OpenStudy (anonymous):

try to solve it atleast

OpenStudy (rayford):

1,2,7,and 14?

OpenStudy (anonymous):

how did u get that ?

OpenStudy (rayford):

honostly...i have no idea.. ._.

OpenStudy (anonymous):

2x^3 – 4x^2 – 7x + 14 = 0 2x^2(x - 2) -7(x - 2) = 0 x=2 or x = square root of 7/2

OpenStudy (rayford):

wait what?? here are the answer choices..

OpenStudy (rayford):

1/2,1,2,7/2,7,14 1,2,7,4....all plus or minus. 1/2,1,2,7/2,7,14 all plus or minus. 1,2,7,14

OpenStudy (rayford):

FOR PEOPLE IN THE FUTURE.....IT'S THE THIRD ANSWER I POSTED.

OpenStudy (rayford):

YOUR WELCOME ;)